The current I (in amps), the power P (in watts), and the resistance R (in ohms) are related by the formula
I = radical over p over r

What current is needed for a 2,700-watt hair dryer if the resistance is 36 ohms? First give the exact answer, expressed as a simplified radical expression.
I = amps

Then give an approximation, rounded to the nearest tenth.
I = amps


Answer by Alan3354(69443)

The current I (in amps), the power P (in watts), and the resistance R (in ohms) are related by the formula
I = radical over p over r
========
Idk what that means.

What current is needed for a 2,700-watt hair dryer if the resistance is 36 ohms? First give the exact answer, expressed as a simplified radical expression.
I = amps
P = I^2 R
2700 = I^2*36
I^2 = 75
I = 5sqrt(3) amps
======================
Then give an approximation, rounded to the nearest tenth.
do that.
I = amps